Operations Report – April 12th

Crew 245 Operations Report 12-04-2021

SOL: 2

Name of person filing report: Shravan Hariharan

Non-nominal systems: Nothing to report

Notes on non-nominal systems: Nothing to report

Generator: Rental generator being installed tonight

Hours run: N/A

From what time last night: N/A

To what time this morning: N/A

List any additional daytime hours when the generator was run: N/A

Solar— SOC 89% (Before generator is run at night)

Diesel Reading – 50%

Station Propane Reading – 75%

Water (loft tank): 31 gallons

Water Meter: 150204.5 units

Water (static tank): 100 gallons

Static to Loft Pump used – yes

Water in GreenHab: 0 gallons

Water in ScienceDome: 0 gallons

Toilet tank emptied: no

Sojourner rover used: ASSIGNED TO DIRECTOR

Hours: N/A

Beginning charge: N/A

Ending charge: N/A

Currently charging: N/A

Spirit rover used: no

Hours: N/A

Beginning charge: N/A

Ending charge: N/A

Currently charging: yes, in town

Opportunity rover used: no

Hours: N/A

Beginning charge: N/A

Ending charge: N/A

Currently charging: no, in town

Curiosity rover used: no

Hours: N/A

Beginning charge: N/A

Ending charge: N/A

Currently charging: yes, in town

Perseverance rover used: Yes

Hours: 199.7

Beginning charge: 100%

Ending charge: 100%

Currently charging: Yes

Notes on rovers: Nothing to report

ATV’s Used: 350.1, 350.2, 350.3

Reason for use: Training

Oil Added? No

ATV Fuel Used: N/A

# Hours the ATVs were used today: 0.5

Notes on ATVs: Mission Director will monitor and refuel ATVs for the duration of the mission.

HabCar used and why, where?: No.

CrewCar used and why, where?: Used on Sol 1 for transport from Grand Junction

CrewCar mileage: 194098

CrewCar fuel: Full

CrewCar oil: 30 quarts, Dipstick full

General notes and comments: Hab kitchen sink is leaking, and seal will be re-caulked around the border of the sink. Leak is minor, and should not require any outside assistance.

Summary of internet: Nominal, crew has not fully consumed daily internet bandwidth so far

Summary of suits and radios: All radios nominal, suits currently charging. Two suit batteries will be swapped tonight, awaiting instructions from Scott Davis.

Summary of Hab operations: Water filter changed and loft tank re-filled, furnace filter changed

Summary of GreenHab operations: None

Summary of ScienceDome operations: Mopped and cleaned by crew

Summary of RAM operations: Inventoried by HSO for safety equipment

Summary of any observatory issues: Nothing to report

Summary of health and safety issues: Nothing to report

Questions, concerns and requests to Mission Support: Nothing to report